Let A=[top(1 1) bottom (0 1)] be a matrix consisting of real numbers (we are not interpreting this as a bit matrix). Find a formula for An, where n ϵ Z+, and use Mathematical Induction to prove that your formula is correct.
1
Expert's answer
2020-12-01T05:47:49-0500

Let "A=\left[\begin{array}{cc} 1 & 1 \\ 0 & 1\end{array}\right]."


Let us prove by Mathematical Induction the following formula:


"A^n=\left[\begin{array}{cc} 1 & n \\ 0 & 1\end{array}\right]."



Base case:


"n=1:"


"A^1=\left[\begin{array}{cc} 1 & 1 \\ 0 & 1\end{array}\right]"

"n=2:"


"A^2=\left[\begin{array}{cc} 1 & 1 \\ 0 & 1\end{array}\right]\cdot \left[\begin{array}{cc} 1 & 1 \\ 0 & 1\end{array}\right]=\left[\begin{array}{cc} 1 & 2 \\ 0 & 1\end{array}\right]"



Inductive step:

Assume that for "n=k" it is true that "A^k=\left[\begin{array}{cc} 1 & k \\ 0 & 1\end{array}\right]" and prove for "n=k+1:"


"A^{k+1}=A^k\cdot A=\left[\begin{array}{cc} 1 & k \\ 0 & 1\end{array}\right]\cdot\left[\begin{array}{cc} 1 & 1 \\ 0 & 1\end{array}\right]=\left[\begin{array}{cc} 1 & k+1 \\ 0 & 1\end{array}\right]"



Conclusion:

Therefore, by Mathematical Induction "A^n=\left[\begin{array}{cc} 1 & n \\ 0 & 1\end{array}\right]" for each "n \in\mathbb Z_+"